OpenStudy (anonymous):

find the limit as x approaches 0 for cos(2x) / x

OpenStudy (anonymous):

(cosxsinx - tanx)/(x^2 sinx) sinx[cosx - (1/cosx)]/(x^2 sinx) [cosx - (1/cosx)]/x^2 (cos^2x - 1) /x^2 cosx - sin^2 x / x^2 cosx = -[sin(x)/x]^2 / cosx limit x-->0 -[sin(x)/x]^2 / cosx = -(1)^2 / 1 = -1 since we know limit x-->0 sin(x) / x = 1 and cos 0 is 1

OpenStudy (anonymous):

O_o

OpenStudy (anonymous):

I think there's an easier way to do this . . .

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

haha. if theres an easier way, please explain.

OpenStudy (anonymous):

Well, cosine of whatever ranges between -1 and 1 right?

OpenStudy (anonymous):

use sandwich theory..

OpenStudy (anonymous):

cos(2x) is varing -1 to 1

OpenStudy (anonymous):

and anything between -1 and 1 divided by zero is undefined, so the limit doesn't exist.

OpenStudy (anonymous):

i took a mistake no sandwich

OpenStudy (anonymous):

I'm going to go eat a sandwich. I'll be back in a couple minutes if you need me.

OpenStudy (anonymous):

ugh. trig functions and limits together bug me so much.

OpenStudy (anonymous):

use L'hospital theory.

OpenStudy (anonymous):

I think L'Hopital's is unreliable here.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Well, it works if you take it far enough, but is more work than is necessary. The reasoning I provided leads to the same conclusion. It's division by zero regardless.

OpenStudy (anonymous):

If you only differentiate once, you'll get -2sin(2x)/1 and if you plug in x=0 now you'll mistakenly believe that the limit is zero. If you keep going and differentiate again, you'll get -4cos(2x)/0 which leads to the same conclusion that I made above.

OpenStudy (anonymous):

why it is possible.

OpenStudy (anonymous):

it can be convetred when the function combined together

OpenStudy (anonymous):

I don't know what you mean.

OpenStudy (anonymous):

and L'hosipital theory is appliable right before the denominator is not zero

OpenStudy (anonymous):

if certain function f(x) when x goes to 0 assume that its value is infinity and g(x) which is denominator goes to 0 but f/g when x goes to zero, it shoul have certain value

OpenStudy (anonymous):

I understand the theory, the limit still does not exist because of division by zero.

OpenStudy (anonymous):

ah.. . i had mistake

OpenStudy (anonymous):

it is not the 0/0 or infinit/0 or ...

OpenStudy (anonymous):

I'm saying that it can be solved with a little logic with no need for calculation. Remember this, "if theres an easier way, please explain."

OpenStudy (anonymous):

it just ... constant/smalllll value

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

so you are right...

OpenStudy (anonymous):

it is just infinity....

OpenStudy (anonymous):

No, it is not infinity; it is undefined.

OpenStudy (anonymous):

......... constant/really small vaue is infinity.. 1/x is ex

OpenStudy (anonymous):

If you approach 0 from the left, the left side limit goes to -∞. Approaching from the right, the right side limit goes to +∞.

OpenStudy (anonymous):

Yeah i meant both of two..

OpenStudy (anonymous):

okay so if i was going to find the limit as x approaches 0 for sin(3x) / x , would that not exist either?

OpenStudy (anonymous):

What if x is a really small negative number? What if x is a really small negative number? Right, so you can't claim that the limit is infinity.

OpenStudy (anonymous):

alright. but i meant that infinity means both of + or -

OpenStudy (anonymous):

if numerator is sine, then you can use L'hostpital. it is 0/0

OpenStudy (anonymous):

That's fine as long as the final answer is "The limit does not exist." The limit for sin(3x)/x does exist.

OpenStudy (anonymous):

im confused. if you graph cos(2x) / x, the limit is infinity isnt it?

OpenStudy (anonymous):

No, if you graph cos(2x)/x the limit at x=0 is undefined, i.e. it does not exist.

OpenStudy (anonymous):

and for sin(3x) / x, it would be 3

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i understand it when i plug it into a graphing calculator. i just dont understand how to do it algebraically.

OpenStudy (anonymous):

LeeYeong's suggestion of L'Hopital's rule is appropriate here.

OpenStudy (anonymous):

One way to think of it is to imagine the triangle that represents the sin(x)/x situation.